OpenStudy (pokiedokie):

x^2 - 10x + 24 = 0 is it x = -4 and x = -6 or is it x = 4 and x = 6 very confused, please help!

ganeshie8 (ganeshie8):

plug \(x = 4\) in the equation and see if it satisfies

OpenStudy (pokiedokie):

okay, give me a minute

OpenStudy (pokiedokie):

it doesn't seem to work

OpenStudy (pokiedokie):

but i factored the equation so it's left to either x = 4 or x = 6, or, x = -4 or x =-6, and i get mixed up with negatives so i guess i just need reassurance

OpenStudy (pokiedokie):

i think it might be the negatives

ganeshie8 (ganeshie8):

x^2 - 10x + 24 = 0 replace x by 4 4^2 - 10*4 + 24 = 0 16 - 40 + 24 = 0 0 = 0 which is a true statement, so 4 is indeed a solution

ganeshie8 (ganeshie8):

it is important that you know how to check if a number satisfies an equation or not

OpenStudy (pokiedokie):

okay, next time i'll try and work through it slower, thank you very much for helping me, i appreciate it!